Problem 6-23(Algorithmic)The medical communityunanimously agrees on the health benefits of regular exercise, but...

  1. Problem 6-23 (Algorithmic)

    The medical community unanimously agrees on the health benefits of regular exercise, but are adults listening? During each of the past 15 years, a polling organization has surveyed americans about their exercise habits. In the most recent of these polls, slightly over half of all American adults reported that they exercise for 30 or more minutes at least three times per week. The following data show the percentages of adults who reported that they exercise for 30 or more minutes at least three times per week during each of the 15 years of this study.


    YearPercentage of Adults Who Exercise 30 or more minutes at least three times per week
    141.5
    245.2
    347.1
    445.6
    546.7
    644.7
    747.9
    850
    948.2
    1049.4
    1150
    1252.4
    1351.1
    1454.9
    1552.5




    1. Choose the correct time series plot.


    Plot (ii)

    Does a linear trend appear to be present?

    Possible Positive Linear Trend

    1. use simple linear regression to find the parameters for the line that minimizes MSE for this time series. Do not round your interim computations and round your final answers to four decimal places. For subtractive or negative numbers use a minus sign. (Example: -300)

      y-intercept, b0 =

      Slope, b1 =

      MSE =

    2. Use the trend equation from part (b) to forecast the percentage of adults next year (year 16 of the study) who will report that they exercise for 30 or more minutes at least three times per week. Do not round your interim computations and round your final answers to four decimal places. For subtractive or negative numbers use a minus sign. (Example: -300)

      %

    3. Use the trend equation from part (b) to forecast the percentage of adults three years from now (year 18 of the study) who will report that they exercise for 30 or more minutes at least three times per week. Do not round your interim computations and round your final answers to four decimal places. For subtractive or negative numbers use a minus sign. (Example: -300)

      %
Answer #1

(a) I am not able to see the plots since the plots did not load in my browser. However, the time series plot would look like the diagram below. I have added the data labels and markers (triangles) for clear visualization of the time series plot.

% of adults who exercise 30 or more minutes at least 3 times per week 52.5 51.1 482 49.4 50 52.4 5 0 47. 9 45.2 47.1 45.6 46.

(b) y-intercept, b0 = 42.7085714285714 = 42.7086

slope, b1 = 0.721428571428571 = 0.7214

Hence, the estimated equation is: % of adults who exercise 30 or more minutes at least 3 times per week = 42.7086 + 0.7214 * Year

MSE = SSE/(n-k) = 26.8954285714286/(15-2) = 2.0689, where n = number of observations and k = number of parameters, SSE = error sum of square

The calculated values are as follows:

6.9 Year Yt Yt - Y_bar X_t-X_bar (Y_t-Y_bar) (X_t - X_bar) (x_t - X_bar)^2Y_hat (Prediction) (residual)^2 1) 41.5 -6.98 -7.00

Steps of estimation as shown below (the formula view of the excel)

A B Year C Y_t b1 SSE =F23/G23 =C22-L4*B22 =SUM(12:116) =L6/(COUNT(B2:B16)-COUNT(L4:L5)) MSE 41.5 45.2 47.1 45.6 46.7 44.7 47

(c) The forecast for year 16 = 42.7085714285714 + 0.721428571428571 * 16 = 54.2514 (refer to the yellow highlighted row above with year = 16)

(d) The forecast for year 16 = 42.7085714285714 + 0.721428571428571 * 18 = 55.6943 (refer to the yellow highlighted row above with year = 18)